# Comparing Specialised Small and General Large Language Models on Text Classification: 100 Labelled Samples to Achieve Break-Even Performance

### v2

@@ Abstract
When solving NLP tasks with limited labelled data, researchers can either use a general large language model without further update, or use a small number of labelled examples to tune a specialised smaller model. In this work, we address the research gap of how many labelled samples are required for the specialised small models to outperform general large models, while taking the performance variance into consideration. By observing the behaviour of fine-tuning, instruction-tuning, prompting and in-context learning on 7 language models, we identify such performance break-even points across 8 representative text classification tasks of varying characteristics. We show that the specialised models often need only few samples (on average $10-1000$) to be on par or better than the general ones. At the same time, the number of required labels strongly depends on the dataset or task characteristics, with this number being significantly lower on multi-class datasets (up to $100$) than on binary datasets (up to $5000$). When performance variance is taken into consideration, the number of required labels increases on average by $100-200\%$ and even up to $1500\%$ in specific cases.
@@ 1 Introduction
With the introduction of the GPT-3 model Brown et al. (2020), large language models have been shown to be an effective generalist models for learning with limited labelled data. They are able to perform well across many NLP tasks with no (using prompting) or only few (using in-context learning) labelled samples and without any parameter update Qin et al. (2023); Sun et al. (2023); Dong et al. (2022). This performance is achieved by conditioning the model on an appropriate text input (prompt) containing instructions for the given task, a test sample for which to generate output and optionally a set of few in-context examples showcasing the task Sun et al. (2023); Dong et al. (2022). Many enhancements were proposed to improve the overall few-shot behaviour. For example prompt-tuning and automatic prompt-engineering, where the effective prompts are designed automatically, or instruction-tuning, where language models are tuned to better follow the task instructions Gao et al. (2021); Logan IV et al. (2022); Dong et al. (2022).
Figure 1: Comparison between the performance of specialised small and general large language models. The break-even points are identified by observing the impact of changing the number of available labelled samples and taking performance variance into consideration. The results are aggregated across the language models relevant for each approach and across 8 datasets (from the individual results presented in Figure 2). Specialised models outperform general ones with only few labelled samples (up to $100$), with performance variance showing strong impact on the comparison, increasing the number significantly.
@@
So far, most of studies comparing the performance of specialised and general language models have been done using different (often non-representative and incomplete) settings and methodologies, leading to divergent findings. The specialised language models are tuned either on the whole labelled set (where they achieve higher performance than general language models) Schick and Schütze (2021); Qin et al. (2023), or using the same set of few samples (e.g., $4-32$) used for general models (where they achieve significantly worse performance than general models) Ma et al. (2023). Only few works study the behaviour of models between the few samples and the whole set, and observe the performance break-even points where specialised language models outperform general ones. Moreover, they often focus only on specific approaches or ignore the performance variance (caused by various sources of randomness, such as random seeds) Le Scao and Rush (2021); Hongjin et al. (2022); Hernandez et al. (2023); Gupta et al. (2023).
@@
Our goal is to remedy these shortcomings and answer the following question: "How many labelled samples do the specialised smaller models need to outperform more general larger models?" To achieve this, we investigate and observe how the performance of different approaches changes when increasing the number of labelled training samples. We identify the break-even points between performance of specialised models and more general models, while taking the performance variance into consideration (aggregated results are presented in Figure 1). Our main contributions and findings are [^1] To support replicability and extension of our results, we openly publish the source code of our experiments at https://anonymous.4open.science/r/investigation-with-interactions-F6D4:
@@
  • We perform a comprehensive investigation of the impact of increasing the number of labelled training samples on performance and its variance of fine-tuning, prompting, in-context learning and instruction-tuning approaches across 7 language models and 8 representative text classification datasets of various characteristics, such as number of classes, text length or task type.
@@
  • By identifying break-even points between the specialised and general models, we find that the smaller specialised models (obtained through fine-tuning or instruction-tuning), require only small number of labelled training examples ($10-1000$) to achieve performance on par or even better than general large language models used in zero/few-shot settings. The required number of labelled samples is dependent on the dataset characteristics, with binary datasets and datasets that require better language understanding (e.g., question answering) requiring significantly more labelled samples for the specialised models to outperform the general ones. In addition, we find significant impact of performance variance, especially originating from in-context learning or fine-tuning on few samples, increasing the number of required labelled samples by up to $1500\%$, with average increase of $100\%-200\%$.
@@ 2 Related Work
Extensive focus is dedicated to the evaluation and comparison between different data efficient approaches utilising large language models, such as prompting, in-context learning, fine-tuning, prompt-tuning or prompt-based/instruction tuning Dong et al. (2022); Liu et al. (2023). The comparisons usually focus on a specific setting and methodologies, such as different model sizes, used approaches or number of labelled samples, and often lead to divergent findings. In some comparisons, the performance of specialised small models is compared with significantly larger general models Schick and Schütze (2021); Ma et al. (2023); Liu et al. (2022); Hernandez et al. (2023); Hongjin et al. (2022), while in others the setting is kept as similar as possible (i.e., comparing specialised and general models of the same sizes) Mosbach et al. (2023); Qin et al. (2023); Logan IV et al. (2022); Gao et al. (2021); Le Scao and Rush (2021); Gupta et al. (2023).
@@
Majority of the comparisons focus on comparing fine-tuned models with in-context learning Ma et al. (2023); Hongjin et al. (2022); Hernandez et al. (2023); Qin et al. (2023), or on comparing prompt-based/instruction-tuned models on the specific tasks with their general counterparts Liu et al. (2022); Mosbach et al. (2023); Schick and Schütze (2021). Only a limited number of papers compare fine-tuning with instruction-tuning Gupta et al. (2023); Le Scao and Rush (2021) or multiple approaches at the same time (i.e., fine-tuning, instruction-tuning, prompting and in-context learning at the same time) Logan IV et al. (2022); Gao et al. (2021).
@@
If enough labelled samples are used for the tuning, the smaller models, specialised either through fine-tuning or instruction-tuning, can achieve performance on par with, or in many cases even better than the performance of the larger general models (using prompting or in-context learning) Schick and Schütze (2021); Qin et al. (2023); Hernandez et al. (2023). At the same time, in the extremely limited settings, where the models are fine-tuned using the same low number of samples as it is used for in-context learning, the general large language models excel over the small specialised models Ma et al. (2023); Hongjin et al. (2022). Other papers study the impact of varying the training data sizes on the performance of specialised models and their comparison with general models. However, they often either perform a comprehensive comparison across multiple approaches, but focus only on few samples and vary the data sizes only to up 32 samples Logan IV et al. (2022); Gao et al. (2021), or focus on small subset of approaches while using a larger part or the whole dataset, but often in specific domains Le Scao and Rush (2021); Gupta et al. (2023); Hongjin et al. (2022); Hernandez et al. (2023); Ma et al. (2023). Only in specific cases the sensitivity of the compared approaches to the sources of randomness (e.g., choice of samples or their order for in-context learning) or to more systematic choices (e.g., choice of prompt format) and the performance variance these factors introduce is taken into consideration Ma et al. (2023).
@@
Compared to these works, we focus on more comprehensive comparisons across: 1) full training dataset, increasing the size from 10 samples to full dataset in exponential fashion; 2) multiple approaches for obtaining specialised models (fine-tuning, prompt-based/instruction-tuning) and for using the general models (prompting, in-context learning); and 3) multiple runs to carefully take into consideration the sensitivity of different approaches and the performance variance this sensitivity introduces.
3 Investigating the Impact of Dataset Size
Approaches.
We investigate the impact of training dataset size across a set of currently popular approaches for dealing with limited labelled data in NLP: 1) fine-tuning; 2) prompting; 3) in-context learning; and 4) instruction-tuning. For fine-tuning, we use the BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) base models. For prompting and in-context learning the Flan-T5 Chung et al. (2024) base, ChatGPT (3.5-turbo-0613 version) Brown et al. (2020); Ouyang et al. (2022), LLaMA-2 Touvron et al. (2023) 13B chat optimised model, Mistral-7B Jiang et al. (2023) and Zephyr-7B Tunstall et al. (2023) models. For instruction-tuning we use the Flan-T5 (full instruction-tuning), Mistral-7B and Zephyr-7B models (both instruction-tuned using LoRA Hu et al. (2021)).

### v3

@@ Abstract
When solving NLP tasks with limited labelled data, researchers typically either use a general large language model without further update, or use a small number of labelled samples to tune a specialised smaller model. In this work, we answer an important question – how many labelled samples are required for the specialised small models to outperform general large models, while taking the performance variance into consideration. By observing the behaviour of fine-tuning, instruction-tuning, prompting and in-context learning on 8 language models, we identify such performance break-even points across 8 representative text classification tasks of varying characteristics. We show that the specialised models often need only few samples (on average $100$) to be on par or better than the general ones. At the same time, the number of required labels strongly depends on the dataset or task characteristics, with fine-tuning on binary datasets requiring significantly more samples. When performance variance is taken into consideration, the number of required labels increases on average by $100-200\%$. Finally, larger models do not consistently lead to better performance and lower variance, with 4-bit quantisation having negligible impact.
@@ 1 Introduction
With the introduction of the GPT-3 model Brown et al. (2020), large language models have been shown to be an effective generalist models for learning with limited labelled data. They are able to perform well across many NLP tasks with no (using prompting) or only few (using in-context learning) labelled samples and without any parameter update Qin et al. (2023); Sun et al. (2023); Liu et al. (2023). This performance is achieved by conditioning the model on an appropriate text input (prompt) containing instructions for the given task, a test sample for which to generate output and optionally a set of few in-context examples showcasing the task Sun et al. (2023); Dong et al. (2022). Many enhancements were proposed to improve the overall few-shot behaviour. For example prompt-tuning and automatic prompt-engineering, where the effective prompts are designed automatically, or instruction-tuning, where language models are tuned to better follow the task instructions Gao et al. (2021); Logan IV et al. (2022).
Figure 1: Comparison between the performance of specialised small and general large language models. The break-even points are identified by observing the impact of changing the number of available labelled samples and taking performance variance into consideration. Specialised models outperform general ones with only few labelled samples (up to $100$), with performance variance showing strong impact on the comparison, increasing the number significantly.
@@
So far, most of the studies comparing the performance of specialised and general language models from the data-efficiency perspective have been done using different (often non-representative and incomplete) settings and methodologies, leading to divergent findings (Pecher et al., 2024b). The specialised language models are tuned either on the whole labelled set (where they achieve higher performance than general language models) Schick and Schütze (2021); Qin et al. (2023), or using the same set of few samples (e.g., $4-32$) used for general models (where they achieve significantly worse performance than general models) Ma et al. (2023). Only few works study the behaviour of models between the few samples and the whole set, and observe the performance break-even points where specialised language models outperform general ones. Moreover, they often focus only on specific approaches or ignore the performance variance (caused by various sources of randomness, such as random seeds) Le Scao and Rush (2021); Hongjin et al. (2022); Hernandez et al. (2023); Gupta et al. (2023); Pecher et al. (2024b).
In addition, almost no work considers the compute-efficiency of the different approaches in the comparisons, even though it is an important aspect when the approaches achieve comparable results (Mosbach et al., 2023). Although some works investigate the impact of quantisation on model performance (Liu et al., 2024; Jin et al., 2024), the impact on the performance variance (i.e., sensitivity) or the overall comparison is not explored.
@@
Our goal is to remedy these shortcomings and answer the following question: "How many labelled samples do the specialised smaller models need to outperform more general larger models?" To achieve this, we investigate and observe how the performance of different approaches changes when increasing the number of labelled training samples (data-efficiency). We identify the break-even points between performance of specialised models and more general models of different sizes (compute-efficiency), while taking the performance variance into consideration (sensitivity to randomness). The aggregated results are presented in Figure 1. Our main contributions and findings are [^1] To support replicability and extension of our results, we openly publish the source code of our experiments at https://anonymous.4open.science/r/investigation-with-interactions-F6D4:
@@
  • We perform a comprehensive and fair investigation of the impact of increasing the number of labelled training samples on performance and its variance of fine-tuning, prompting, in-context learning and instruction-tuning approaches across 8 language models and 8 representative text classification datasets of various characteristics, such as number of classes, text length or task type.
@@
  • By identifying break-even points between the specialised and general models, we find that the smaller specialised models (obtained through fine-tuning or instruction-tuning), require only a small number of labelled training examples ($100$ on average) to achieve performance on par or even better than general large language models used in zero/few-shot settings. In addition, we find significant impact of performance variance, especially originating from in-context learning or fine-tuning on few samples, increasing the number of required labelled samples by up to $500-3000\%$, with average increase of $100\%-200\%$.
  • Based on further analysis we find following key insights: 1) the required number of labelled samples is dependent on the dataset characteristics, with binary datasets and datasets that require better language understanding (e.g., question answering) requiring more labelled samples; 2) larger models do not necessarily lead to better model performance, data-efficiency or lower performance variance for general models; and 3) 4-bit quantisation has negligible impact on the overall behaviour of general models
@@ 2 Related Work
Extensive focus is dedicated to the evaluation and comparison between different data efficient approaches utilising large language models, such as prompting, in-context learning, fine-tuning, prompt-tuning or prompt-based/instruction tuning on text classification tasks Dong et al. (2022); Liu et al. (2023). In some comparisons, the performance of specialised small models is compared with significantly larger general models Schick and Schütze (2021); Ma et al. (2023); Liu et al. (2022); Hernandez et al. (2023); Hongjin et al. (2022), while in others the setting is kept as similar as possible (i.e., comparing specialised and general models of the same sizes) Mosbach et al. (2023); Qin et al. (2023); Logan IV et al. (2022); Gao et al. (2021); Le Scao and Rush (2021); Gupta et al. (2023).
@@
Majority of the comparisons focus on comparing fine-tuned models with in-context learning Ma et al. (2023); Hongjin et al. (2022); Hernandez et al. (2023); Qin et al. (2023), or on comparing prompt-based/instruction-tuned models on the specific tasks with their general counterparts Liu et al. (2022); Mosbach et al. (2023); Schick and Schütze (2021). Only few papers compare fine-tuning with instruction-tuning Gupta et al. (2023); Le Scao and Rush (2021) or multiple approaches at the same time (i.e., fine-tuning, instruction-tuning, prompting and in-context learning at the same time) Logan IV et al. (2022); Gao et al. (2021).
@@
If enough labelled samples are used, the smaller specialised models can achieve performance on par with, or in many cases even better than the performance of the larger general models Schick and Schütze (2021); Qin et al. (2023); Hernandez et al. (2023). At the same time, in the extremely limited settings, where the models are fine-tuned using the same low number of samples as it is used for in-context learning, the general large language models excel over the small specialised models Ma et al. (2023); Hongjin et al. (2022). Other papers study the impact of varying the training data sizes on the performance of specialised models and their comparison with general models. However, they often either perform a comprehensive comparison across multiple approaches, but focus only on few samples and vary the data sizes only to up 32 samples Logan IV et al. (2022); Gao et al. (2021), or focus on small subset of approaches while using a larger part or the whole dataset, but often in specific domains Le Scao and Rush (2021); Gupta et al. (2023); Hongjin et al. (2022); Hernandez et al. (2023); Ma et al. (2023). Only few papers consider (to a certain extent) effects of different systematic choices (e.g., number of samples, or prompt format) or sensitivity to sources of randomness (e.g., choice or order of samples) on comparison (Ma et al., 2023; Pecher et al., 2024a; Sclar et al., 2023; Weber et al., 2023).
@@
Compared to these works, we focus on more comprehensive comparisons across: 1) full training dataset, increasing the size from 10 samples to full dataset in exponential fashion; 2) multiple approaches for obtaining specialised models (fine-tuning, prompt-based/instruction-tuning) and for using the general models (prompting, in-context learning); 3) multiple runs to carefully take into consideration the sensitivity of different approaches and the performance variance this sensitivity introduces; and 4) models of different sizes.
3 Comparison Methodology: Identifying Performance Break-Even Points
Our main focus is on comparing the specialised and general models from the perspective of their data- and compute-efficiency and sensitivity to the effects of randomness, as illustrated by Figure 1.
Data-efficiency.
To compare the approaches from the perspective of the data-efficiency, we observe and compare how their performance changes when increasing the number of available labelled samples from low number of samples up until the full dataset is used (where each approach is presented with the same set of samples). As we change the number of available samples, we identify the first break-even point between the performance of the specialised and general models. The first break-even point (average performance) specifies the point after which the performance of the specialised models is better on average, but may still be lower in many cases due to the performance variance and the randomness sensitivity, such as when the randomness is not sufficiently addressed (low number of runs is used or the runs are cherry-picked). At each point, we report the mean F1 macro and standard deviation.
Compute-efficiency.
We compare models of different sizes (e.g., using significantly smaller specialised models) while specifically focusing on models that require a comparable amount of computation. For example, full training and evaluation of the smallest model should require the same amount of compute as using the smallest LLM through prompting. More information are provided in Appendix A.3. In addition, for selected general models, we explore the impact of quantisation on the performance and its variance.
Sensitivity to the effects of randomness.
To further explore the impact of the sensitivity to the effects of randomness on the comparison, we identify the second break-even points (average+variance). This break-even point denotes the "worst-case size" or the point after which the specialised models show better performance even when the variance is taken into consideration. In other words, there is low probability that the performance of the specialised models will be lower even when using low number of runs or cherry-picking them. To identify this break-even point, we repeat the training and evaluation multiple times and then compare the worst case performance of the better performing model (obtained by subtracting the standard deviation from the mean) with the best case performance of the worse performing model (obtained by adding the standard deviation to the mean). Each repeat uses the same fixed random seed for all models to guarantee deterministic and replicable results. The random seed covers all sources of randomness (Gundersen et al., 2022a; Pecher et al., 2024b) – the split of data, selection of labelled samples, initialisation of models, order of data in training, non-deterministic operations (e.g., dropout), and selection of samples for in-context learning.
@@ 3 Comparison Methodology: Identifying Performance Break-Even Points
Approaches and models.
We investigate the impact of training dataset size across a set of currently popular approaches for dealing with limited labelled data in NLP: 1) fine-tuning; 2) instruction-tuning; 3) prompting; and 4) in-context learning. For fine-tuning, we use the BERT Devlin et al. (2019a) and RoBERTa Liu et al. (2019) base models. For prompting and in-context learning the Flan-T5 Chung et al. (2024) base, GPT (4o-mini-2024-07-18 version) Brown et al. (2020); Ouyang et al. (2022), LLaMA-2 Touvron et al. (2023) 13B chat optimised model, LLaMA-3 (Dubey et al., 2024) 8B instruction optimised model, Mistral-7B Jiang et al. (2023) and Zephyr-7B Tunstall et al. (2023) models. For instruction-tuning we use the Flan-T5 (full instruction-tuning), Mistral-7B and Zephyr-7B models (both instruction-tuned using LoRA Hu et al. (2021)). Due to the significant costs, we use 4-bit quantisation for the LLaMA-2 model. In case of Mistral-7B and Zephyr-7B, we use both the full-precision and the 4-bit quantised versions, reporting only the best performance.